# Less is more: Probabilistic reduction is best explained by small-scale predictability measures

### v1

@@ 5 N-gram versus language model predictability and probabilistic reduction in upadhye2025back
Probabilistic reduction may not apply as consistently across all words in the lexicon. Specifically, function words typically play a more structural or morphosyntactic role in English, and content words are more interchangeable. Consequently, some have claimed that function words are retrieved and planned at an earlier stage than content words garrett1975analysis. Consistent with this, bell2009predictability describe a dissociation in their analysis of Switchboard, in which content words were only reduced when backward transition probabilities were high, but function words were only reduced when forward transition probabilities were high.
@@
An alternative position to the separate planning of content and function words is that the content-function distinction is somewhat fraught or arbitrary, or mostly reflects lexical frequency staub2024function. Similarly, if speakers retrieve phrase-shaped representations for production, then it is also plausible that probabilistic reduction should apply to both types of words equally. Such a finding would be consistent with upadhye2025back, whose results contradicted bell2009predictability, such that forward and backward predictability led to reduction on both content and function words in the Switchboard-NXT corpus. [^2] We wish to point out that the content-function distinction in the corpus is somewhat unclear. bell2009predictability state, ”The function-word category included discourse markers, conjunctions, existential there, pronouns and other proforms, prepositions, articles, quantifiers, demonstratives, verb auxiliaries, and verb particles. The remainder, nouns, verbs, adverbs, and adjectives, made up the content-word category.” However, an examination of the dataset reveals many words that belong to both categories, a large number of content words in the function word category, and pronouns notably included in the content word category. Future iterations of work in this vein should clarify or correct the content/function distinction. Therefore, in this experiment we replicate their analyses of reduction on content and function word durations.
@@
The Switchboard-NXT data from upadhye2025back contains 650,570 observations. For language model probabilities, we use the provided autoregressive predictors from their custom GPT-2 model radford_language_2019 trained in both forward and backward directions on the CANDOR corpus333upadhye2025back present many other technological solutions that we find interesting, but these are not the focus of the present work.. For analyses of n-gram fluency, we rely on their smoothed bigram estimates derived from the CANDOR corpus, which serve a similar role to the bigram predictors used in bell2009predictability and seyfarth2014word. We broadly replicate all of their results with bigram statistics, such that we observe probabilistic reduction on both content and function words in both forward and backward directions (Table 3). Importantly, we also find that language model probabilities provide a worse fit to the duration data than bigram probabilities ($\Delta LL=2562$, $\chi^{2}=5123.6$, $p$ < .001).
@@
One concern about these results (raised in Footnote 2) is that the presence of reduction for both Content and Function words could be partly due to errors in the coding of Content or Function word category labels, or due to the somewhat arbitrary distinction between the two. We thus conducted an additional analysis without the word class distinction and interaction terms and obtained roughly similar results, again finding that probabilistic reduction in both forward and backward directions is observable with simple bigram statistics. However, future work should determine the extent to which label noise negatively impacts statistical power to detect differences in patterns of probabilistic reduction across content and function words.
@@
(Intercept) 0.138 0.022 6.096 <.001
SUBTLEX word frequency log -0.124 0.004 -32.227 <.001
Number of phonemes 0.133 0.002 86.337 <.001
Utterance length in words ($n$) log 0.008 0.001 7.047 <.001
Function words log -0.211 0.012 -17.076 < .001
Forward bigram probability (Content) log -0.007 0.001 -38.126 <.001
Backward bigram probability (Content) log -0.034 0.001 -52.036 <.001
Forward bigram probability (Function) log -0.004 0.001 -4.835 <.001
Backward bigram probability (Function) log -0.007 0.001 -7.833 <.001
Relative position inverse log 0.092 0.001 144.307 <.001
@@
One reason for the poor explanatory power of LLM probabilities might lie in an apparent confound between language model estimates and position within an utterance. We visualize several complex relationships between position, length, and duration in Figure 3. For language model probabilities, forward transition probabilities decrease substantially over positions, with the specific trajectory affected slightly by utterance length. Backward predictability furthermore shows a complex, highly non-linear pattern in which early, non-initial words and utterance-final words are assigned greater probabilities than utterance-medial words. Since as words approach the ends of the sentence, durations also increase, the predictive power of language model probabilities may be an artifact of a word’s position within an utterance. Whereas autoregressive language models may encode word position in a sentence more explicitly haviv-etal-2022-transformer, such confounds may be less present with the n-gram estimates as they do not explicitly encode position information, a possibility we explore in Section §6.
@@
Thus, this analysis shows that not only do large language models lack both cognitive plausibility, relative to phrase-based representations that support successful language production arnon2013more; siyanova2018production, they also provide worse fits to the data, possibly due to major confounding factors.
@@
Figure 3: Relationship utterance length and relative position on language model probabilities and duration in Switchboard corpus.
@@
6 Examining probabilistic reduction in smaller and diverse dialect corpora
@@
Previous work on probabilistic reduction has focused on White Mainstream American English (WMAE), such as that spoken in the Buckeye Corpus pitt2005buckeye and Switchboard godfrey1992switchboard; calhoun2010nxt. This work examines both WMAE and several varieties of African American Language (AAL), as AAL varieties are under-represented in computational psycholinguistics (as elsewhere in natural language processing; 10887767). Dialects represent a critical test case for modeling probabilistic language processing, because speech communities, and thus the linguistic statistics of those communities, differ along many dimensions.
@@
A major challenge for assessing probabilistic reduction in dialect corpora, and smaller datasets more broadly, is that large language models may not produce high-quality estimates, either when trained on the corpora directly, or when fine-tuned. Moreover, while there is no reason to expect that predictability would not shape production in a similar way across dialects, if dialects are modeled by variables extracted from mainstream varieties (e.g., the models used in clark_relationship_2025 and upadhye2025back), then these biases are expected to distort any effects of linguistic statistical knowledge on production. This experiment therefore explores the feasibility of applying n-gram statistics to probabilistic reduction across smaller and diverse dialect corpora.
@@
It is important to level the playing field so that LLM probabilities can capture dialect-specific variability in a low-data training regime, particularly since n-gram models trained over a specific corpus necessarily capture variability intrinsic to a dialect. vskrjanec2026language showed that fine-tuning language models (rather than training a small-scale neural language model from scratch) provides one means to capture the knowledge of a person’s specific linguistic experience to explain what they will find difficult to process (or produce). Thus, our experiments in this section compare fine-tuned language models against n-gram models of words’ forward and backward predictability.
6.1 Data preprocessing
@@
In our experiments, we compare and contrast two datasets of spontaneous speech corpora representing Mainstream American English and African American Language. For MAE, we use the Buckeye Corpus used in seyfarth2014word, which contains 282,565 tokens. We compare probabilistic reduction in Buckeye with CORAAL kendall2018corpus. The full CORAAL corpus is much larger than Buckeye, with 1.37 million tokens, and as of this writing is composed of eight distinct dialect groups from across the United States. Since only some dialects had time-aligned transcripts, we applied a CORAAL-specific language model in the Montreal Forced Aligner (MFA) to obtain durations for all recordings mcauliffe2017montreal [^4] The MFA language model was trained on only four dialects: DCA, DCB, PRV, and ROC..
@@
As in Section §4, we tokenized the spoken corpora into utterances along pause boundaries, as these roughly correspond to intonational units bishop2018anticipatory. Buckeye and CORAAL defined pauses somewhat differently, with CORAAL using phonetic criteria (60-70 ms of silence; kendallthesis09; kendall2013speech) and Buckeye relying on the intuition of the annotators pitt2005buckeye; raymond2003analysis. Furthermore, AAL has many grammatical and lexical features that are relatively unfamiliar to the base model, relative to WMAE. So, for the CORAAL data, we merged each utterance with the next to augment the input size and ensure a strong baseline for both datasets.
6.2 Language model fine-tuning
@@
As language model probabilities extracted over short sequences appear to be better predictors of word durations (see Section §4), we chose to fine-tune the Pythia-160M language model biderman2023emergent over inter-pause units for both Buckeye and CORAAL experiments. Pythia is an autoregressive language model with a high degree of psychometric predictive power, and it has been successfully applied to other psychophysical data such as human reading times oh2023transformer; oh2024frequency. Fine-tuning proceeded for both Buckeye and CORAAL models for 40 epochs. We verified convergence by visually inspecting validation set loss using Tensorboard tensorflow2015-whitepaper. We share additional fine-tuning parameters in the Appendix.
6.3 Reverse language model fine-tuning
@@
To obtain an estimate of a word’s backward transition probability, we again start with Pythia-160M as a baseline model, and we separately fine-tune this model on reversed MAE and AAL corpora. We used the same fine-tuning parameters as for the forward-trained models.
@@
The correlation between neural language model probabilities and relative position persists in these data. We visualize the relative lack of correlation, characterized primarily by confounds with utterance edges, that is obtained by a simple trigram model over the corpora in Figure 4. Specifically, utterance-initial words tend to be relatively improbable in trigram models. Forward probabilities typically increase in their predictability after one or two words to a consistent baseline. Backward probabilities similarly are relatively probable for the early words in an utterance, but these increase to relatively high probabilities by the beginning.
@@
Figure 5: Regression model coefficient estimates for probabilistic reduction by corpus, language model type, and direction.
@@
(Intercept) 0.562 0.019 29.24 <.001
SUBTLEX word frequency log -0.132 0.003 -38.29 <.001
Number of phonemes 0.120 0.001 85.36 <.001
Utterance length in words ($n$) log -0.114 0.001 -117.33 <.001
Forward trigram probability log -0.021 0.0002 -67.37 <.001
Backward trigram probability log -0.004 0.0003 -15.72 <.001
Relative position inverse log 0.127 0.001 226.74 <.001
Table 4: Results of CORAAL analysis.
@@
(Intercept) 0.811 0.029 28.30 <.001
SUBTLEX word frequency log -0.190 0.005 -35.99 <.001
Number of phonemes 0.129 0.002 59.42 <.001
Utterance length in words ($n$) log -0.100 0.002 -48.99 <.001
Forward trigram probability log -0.020 0.001 -31.62 <.001
Backward trigram probability log -0.007 0.001 -11.49 <.001
Relative position inverse log 0.121 0.001 95.30 <.001
Table 5: Results of Buckeye analysis.
We attempted to determine whether the different dialect groups in CORAAL varied in the extent to which probabilistic reduction was evident. A model that included dialect did not converge, and reduction effects were stable across all dialects. We visualize the stability of this pattern in Figure 6.
Figure 6: All dialects show the same effect of predictability on duration.
@@ 7 Effect size of probabilistic reduction
As we report, many previous studies have reported contradictory results about the relationship between a word’s predictability and its duration, including our own comparison between LLM and n-gram based predictors. In this section, we illustrate how the coefficient estimates of probabilistic reduction are dwarfed by the estimates of other factors, even though all probabilistic predictors are highly significant and model comparison strongly supports inclusion of these predictors in the full model.
One plausible explanation for inconsistencies across studies of probabilistic reduction effects likely lies in the absolute size of the effect. When considering coefficient $t$ values as a proxy for standardized effect sizes, many predictors, particularly “garden variety” prosodic ones such as the relative position of a word in an utterance, can be an order of magnitude larger than these contextual probability estimates. Figure 7 directly compares the results of the Switchboard re-analysis from Section §5 to the results of the CORAAL and Buckeye analysis.
While all n-gram analyses are consistent with each other in the sense that probabilistic reduction is observed across all experiments, the effects themselves could be easy to miss, and studies may be underpowered, particularly depending on regression model parameterization. This potential explanation is corroborated by the finding that LLM probabilities fail to consistently produce negative coefficients.
@@
[…]
8 Discussion
In this work, we demonstrated that phrase-based representations provide a strong baseline to identify probabilistic reduction in duration. Despite claims by clark_relationship_2025 and upadhye2025back that the power of large language models enables them to uniquely capture probabilistic reduction in forward and backward directions and the effect of givenness or discourse status, we find that such claims are emphatically not supported by the data. That is, once prosodic timing variables specific to whole utterances are considered, particularly word position and utterance length, LLM-based predictors provide worse fits to duration data, potentially due to a number of confounds between position, length, and language model probability estimates. Throughout all of our experiments, it is clear that less context produces higher-quality estimates of speech production data, and more consistently.
Such findings provide strong evidence, counter to the zeitgeist, that small-scale, relatively incremental production, is critical for capturing phonological encoding effects. We argue that the consistent explanatory power of phrase statistics, and smaller-scale large language model probability estimates, suggests that if speakers are in fact using some estimate of a word’s probability in context during the phonological encoding process, they are doing so in a relatively local way. What we observe as effects of rational production that reduce words that are highly predictable may instead reflect other processes cohenpriva26. For example, speakers may retrieve phrases and prosodic knowledge about those phrases from memory somewhat incrementally during production, which could lead to reduction that manifests as a function of probabilities, but which instead reflects something like bigram or trigram frequency bannard2008stored; arnon2013more.
In addition to being poor explainers of phonetic duration, the application of LLMs to the study of probabilistic reduction in general poses several ethical challenges. First, many language models are typically taken "off-the-shelf", without regard to the nature of the web origins of their training data. In addition to domain differences between spontaneous speech and written language, LLMs that superimpose the statistics of the dominant variety on marginalized varieties will ultimately perpetuate hegemonic data practices birhane2021impossibility; birhane2022values. Fine-tuning or training from scratch may serve as one mitigation strategy to reduce the biases caused by these decisions, but clearly this is not necessary or sufficient to show probabilistic reduction.

### v2

@@ 4 Study 2: N-gram versus language model predictability and probabilistic reduction in upadhye2025back
Probabilistic reduction may not apply as consistently across all words in the lexicon. Specifically, function words typically play a more structural or morphosyntactic role in English, and content words are more interchangeable. Consequently, some have claimed that function words are retrieved and planned at an earlier stage than content words (garrett1975analysis). Consistent with this, bell2009predictability describe a dissociation in their analysis of Switchboard, in which content words were only reduced when backward transition probabilities were high, but function words were only reduced when forward transition probabilities were high.
@@
An alternative position to the separate planning of content and function words is that the content-function distinction is somewhat fraught or arbitrary, or mostly reflects lexical frequency (staub2024function). Similarly, if speakers retrieve phrase-shaped representations for production, then it is also plausible that probabilistic reduction should apply to both types of words equally. Such a finding would be consistent with upadhye2025back, whose results contradicted bell2009predictability, such that forward and backward predictability led to reduction on both content and function words in the Switchboard-NXT corpus. [^1] We wish to point out that the content-function distinction in the corpus is somewhat unclear. bell2009predictability state, “The function-word category included discourse markers, conjunctions, existential there, pronouns and other proforms, prepositions, articles, quantifiers, demonstratives, verb auxiliaries, and verb particles. The remainder, nouns, verbs, adverbs, and adjectives, made up the content-word category.” However, an examination of the dataset reveals many words that belong to both categories, a large number of content words are found in the function word category, and pronouns are notably included in the content word category. Future iterations of work in this vein should clarify or correct the content/function distinction. Therefore, in this experiment we replicate their analyses of reduction on content and function word durations.
@@
The Switchboard-NXT data from upadhye2025back contains 650,570 observations. For language model probabilities, we use the provided autoregressive predictors from their custom GPT-2 model (radford_language_2019) trained in both forward and backward directions on the CANDOR corpus222upadhye2025back present many other technological solutions that we find interesting, but these are not the focus of the present work.. For analyses of n-gram fluency, we rely on their smoothed bigram estimates derived from the CANDOR corpus, which serve a similar role to the bigram predictors used in bell2009predictability and seyfarth2014word. We broadly replicate all of the results of upadhye2025back with bigram statistics, such that we observe probabilistic reduction on both content and function words in both forward and backward directions (Table 3). Importantly, we also find that language model probabilities provide a worse fit to the duration data than bigram probabilities ($\Delta LL=2562$, $\chi^{2}=5123.6$, $p$ ¡ .001).
@@
One concern about these results (raised in Footnote 1) is that the presence of reduction for both Content and Function words could be partly due to errors in the coding of Content or Function word category labels, or due to the somewhat arbitrary distinction between the two. We thus conducted an additional analysis without the word class distinction and interaction terms and obtained roughly similar results, again finding that probabilistic reduction in both forward and backward directions is observable with simple bigram statistics. However, future work should determine the extent to which label noise negatively impacts statistical power to detect differences in patterns of probabilistic reduction across content and function words.
@@
(Intercept) 0.138 0.022 6.096 ¡.001
SUBTLEX word frequency log -0.124 0.004 -32.227 ¡.001
Number of phonemes 0.133 0.002 86.337 ¡.001
Utterance length in words ($n$) log 0.008 0.001 7.047 ¡.001
Function words log -0.211 0.012 -17.076 ¡ .001
Forward
Content log -0.007 0.001 -38.126 ¡.001
Function log -0.004 0.001 -4.835 ¡.001
Backward
Content log -0.034 0.001 -52.036 ¡.001
Function log -0.007 0.001 -7.833 ¡.001
Relative position inverse log 0.092 0.001 144.307 ¡.001
@@
One reason for the poor explanatory power of LLM probabilities might lie in an apparent confound between language model estimates and position within an utterance. We visualize several complex relationships between position, length, and duration in the Switchboard data in Figure 3. For language model probabilities, forward transition probabilities decrease substantially over positions, with the specific trajectory affected slightly by utterance length. Backward predictability furthermore shows a complex, highly non-linear pattern in which early, non-initial words and utterance-final words are assigned greater probabilities than utterance-medial words. Because the durations of words toward the ends of the sentence also increase (see Figure 1), the predictive power of language model probabilities relative to n-grams in upadhye2025back may be an artifact of a word’s position within an utterance. Whereas autoregressive language models may encode word position in a sentence more explicitly (haviv-etal-2022-transformer), such confounds may be less present with the n-gram estimates as they do not explicitly encode position information, a possibility we explore in Section §5.
@@
In sum, this analysis shows that not only do large language models lack cognitive plausibility, relative to phrase-based representations that support successful language production (arnon2013more; siyanova2018production), they also provide worse fits to the data, possibly due to major confounding factors.
@@
Figure 3: Relationship between utterance length and relative position on language model probabilities and duration in upadhye2025back
Interim Discussion
Our re-analysis of clark_relationship_2025 and upadhye2025back suggests that positional information very strongly shapes prosodic structure. More importantly, it is clear that prosodic structure is partly confounded with probabilities from large language models, and that longer-range estimates of a word’s predictability do not provide additional explanatory power for word durations. Surprisingly, probabilistic reduction is evident in both datasets, suggesting that a word’s predictability affects its phonological encoding (aylett2004smooth). One major open question is about the generality of probabilistic reduction, both in terms of speech communities and the domain of spontaneous speech under analysis. The final study explores these questions.
@@
5 Study 3: Examining probabilistic reduction in smaller and diverse dialect corpora
The Switchboard and CANDOR corpus are both corpora containing dialogues between strangers who may be likely to vary word durations to accommodate their addressees. Spontaneous speech exists in a variety of domains, however, and speakers’ own pronunciations might change because they are addressing strangers. Interview settings have demonstrated probabilistic reduction: seyfarth2014word showed that words that are more predictable in the Buckeye corpus are produced with shorter durations based on bigram statistics. We specifically test whether the same forward and backward predictability-driven reduction relationships are still observed in more casual, self-driven interviews without the pressure of conversational demands. This final analysis replicates seyfarth2014word and extends his analysis to a larger, more diverse corpus that enables us to ask how universal probabilistic reduction is in spontaneous monologue.
@@ 5 Study 3: Examining probabilistic reduction in smaller and diverse dialect corpora
Previous work on probabilistic reduction has focused on White Mainstream American English (WMAE), such as that spoken in the Buckeye Corpus (pitt2005buckeye) and Switchboard (godfrey1992switchboard; calhoun2010nxt). This study examines both WMAE and several varieties of African American Language (AAL), as AAL varieties are under-represented in computational psycholinguistics (as elsewhere in natural language processing; 10887767). Dialects represent a critical test case for modeling probabilistic language processing, because speech communities, and thus the linguistic statistics of those communities, differ along many dimensions.
@@
A major challenge for assessing probabilistic reduction in dialect corpora, and smaller datasets more broadly, is that large language models may not produce high-quality estimates of contextual probabilities, either when using off-the-shelf models, when trained on the corpora directly, or when fine-tuned. Moreover, while there is no reason to expect that predictability would not shape production in a similar way across dialects, if dialects are modeled by variables extracted from mainstream varieties (e.g., the models used in clark_relationship_2025 and upadhye2025back), then these biases are expected to distort any effects of linguistic statistical knowledge on production. This experiment therefore explores the feasibility of applying n-gram statistics to probabilistic reduction across smaller and more diverse dialect corpora.
@@
It is important to level the playing field so that LLM probabilities can capture dialect-specific variability in a low-data training regime, particularly since n-gram models estimated over a specific corpus necessarily capture variability intrinsic to a dialect. vskrjanec2026language showed that fine-tuning language models provides one means to capture the knowledge of a person’s specific linguistic experience to explain what they will find difficult to process (or produce) as an alternative to training a small-scale neural language model from scratch. Thus, our experiments in this section compare fine-tuned language models against n-gram models of words’ forward and backward predictability.
5.1 Dialect corpus preprocessing
@@
In our experiments, we compare and contrast two datasets of spontaneous speech corpora representing Mainstream American English and African American Language. For MAE, we use the Buckeye Corpus used in seyfarth2014word, which contains 282,565 tokens. We compare probabilistic reduction in Buckeye with CORAAL, which is a broad-coverage corpus of African-American Language (kendall2018corpus). Speakers from this corpus come from Atlanta, Georgia; Washington DC at two different time points; Detroit, Michigan; New York City’s Lower East Side, Princeville, North Carolina; Rochester, New York; and Valdosta, Georgia and thus represent a range of histories of migration and language contact. The full CORAAL corpus is much larger than Buckeye, with 1.37 million tokens, and as of this writing is composed of eight distinct dialect groups from across the United States.
@@
Both corpora clearly demarcate phrase boundaries in the transcripts. As in Section §3, we tokenized the spoken corpora into utterances along pause boundaries, as these roughly correspond to intonational units (bishop2018anticipatory). Buckeye and CORAAL defined pauses somewhat differently, with CORAAL using phonetic criteria (60-70 ms of silence; kendallthesis09; kendall2013speech) and Buckeye relying on the intuition of the annotators (pitt2005buckeye; raymond2003analysis). We suspect that these differences are not major, however, as we showed in Figure 1 that both corpora demonstrate the expected anticipatory shortening and phrase-final lengthening under our segmentation scheme.
We took gold inter-pause intervals and then produced timestamps for word boundaries for all of the recordings in both corpora. Since only some dialects had time-aligned transcripts for word boundaries, for consistency we applied a CORAAL-specific language model in the Montreal Forced Aligner (MFA) to obtain durations of words for all recordings (mcauliffe2017montreal) [^3] The MFA language model was trained on only four dialects: DCA, DCB, PRV, and ROC..
5.1.1 Language model fine-tuning
@@
As language model probabilities extracted over short sequences appear to be better predictors of word durations (see Section §3), we chose to fine-tune the Pythia-160M language model (biderman2023emergent) over inter-pause units for both Buckeye and CORAAL experiments using the zeldarose python package (grobol2023zelda). Pythia is an autoregressive language model with a high degree of psychometric predictive power, and it has been successfully applied to other psychophysical data such as human reading times (oh2023transformer; oh2024frequency). We verified convergence by visually inspecting validation set loss using Tensorboard (tensorflow2015-whitepaper). We share fine-tuning parameters in the Appendix.
In our fine-tuning procedure, we took into account that AAL has many grammatical and lexical features that are relatively unfamiliar to Pythia, relative to WMAE (dayton1996grammatical; mufwene2021african). So, for fine-tuning a language model on the CORAAL data only, we merged each utterance with the next to augment the input size and ensure a strong language modeling baseline for both datasets.
5.1.2 Reverse language model fine-tuning
@@
To obtain an estimate of a word’s backward transition probability, we again start with Pythia-160M as a baseline model, but we reversed the order of all words in each utterance and separately fine-tuned this model on reversed MAE and AAL corpora. We used the same fine-tuning parameters as for the forward-trained models.
5.2 Results
@@
The correlation that we observed in the upadhye2025back data between neural language model probabilities and relative position persists in these data. However, for trigram-based estimates of a word’s probability in context, we observe a relative lack of correlation, characterized primarily by extreme probabilities at utterance edges, which we illustrate in Figure 4. Utterance-initial words tend to be relatively improbable in trigram models. Forward probabilities typically increase in their predictability after one or two words to a consistent baseline. Backward probabilities similarly are relatively probable for the early words in an utterance, but these increase to relatively low probabilities by the end of the sentence.
@@
Figure 5: Regression model coefficient estimates for probabilistic reduction by corpus, language model type, and direction for Study 3.
@@
(Intercept) 0.562 0.019 29.24 ¡.001
SUBTLEX word frequency log -0.132 0.003 -38.29 ¡.001
Number of phonemes 0.120 0.001 85.36 ¡.001
Utterance length in words ($n$) log -0.114 0.001 -117.33 ¡.001
Forward transition probability log -0.021 0.0002 -67.37 ¡.001
Backward transition probability log -0.004 0.0003 -15.72 ¡.001
Relative position inverse log 0.127 0.001 226.74 ¡.001
Table 4: Results of CORAAL analysis with trigram predictability measures.
@@
(Intercept) 0.811 0.029 28.30 ¡.001
SUBTLEX word frequency log -0.190 0.005 -35.99 ¡.001
Number of phonemes 0.129 0.002 59.42 ¡.001
Utterance length in words ($n$) log -0.100 0.002 -48.99 ¡.001
Forward transition probability log -0.020 0.001 -31.62 ¡.001
Backward transition probability log -0.007 0.001 -11.49 ¡.001
Relative position inverse log 0.121 0.001 95.30 ¡.001
Table 5: Results of Buckeye analysis with trigram predictability measures.
We also attempted to determine whether the different dialect groups in CORAAL varied in the extent to which probabilistic reduction was evident. A model that included dialect did not converge, and reduction effects were stable across all dialects. We visualize the stability of this pattern in Figure 6.
Figure 6: All CORAAL dialects show the same effect of predictability on duration.
@@ 6 Effect size of probabilistic reduction
As we report, many previous studies have reported contradictory results about the relationship between a word’s predictability and its duration, including our own comparison between LLM and n-gram based predictors. One plausible explanation for inconsistencies across studies of probabilistic reduction effects likely lies in the absolute size of the effect. This potential explanation is corroborated by the finding that LLM probabilities fail to consistently produce negative coefficients. In this section, we illustrate how the coefficient estimates of probabilistic reduction are dwarfed by the estimates of other factors, even though all probabilistic predictors are highly significant and model comparison strongly supports inclusion of these predictors in the full model.
When considering coefficient $t$ values as a proxy for standardized effect sizes, many predictors, particularly “garden variety” prosodic ones such as the relative position of a word in an utterance, can be an order of magnitude larger than these contextual probability estimates. While all n-gram analyses are consistent with each other in the sense that probabilistic reduction is observed across all experiments, the effects themselves could be easy to miss, and studies may be underpowered, particularly depending on regression model parameterization. Figure 7 illustrates the results of the Switchboard re-analysis from Section §4 alongside the results of the CORAAL and Buckeye analysis from Section §5.
@@
Inconsistency in results is also likely obscured due to changes in statistical reporting practices. It is now commonplace in computational psycholinguistics to justify the inclusion of probabilistic predictors through model comparison, following early work by bell2009predictability. However, one shortcoming of this approach is that the field’s preferred measure, the difference in model log-likelihood ($\Delta LL$), does not include sign information, nor does it transparently convey the absolute value of the coefficient either. That is, it is possible to report a major improvement in model fit that does not necessarily support claims of probabilistic reduction, and such an effect may be quite small. Reliance on $\Delta LL$ likely exacerbates the difficulty of finding a significant true effect of probability on phonetic duration, which appears to be highly sensitive to regression model configuration and conditions.